# Cup and cap products for cohomology and homology groups of ample groupoids

### v1

@@ 1 Introduction
We aim to introduce the cup and cap products in a manner that is both thorough and accessible. Our goal is to ensure that even those who may not be familiar with algebraic topology can easily grasp these concepts. We will carefully explain the definitions and discuss the fundamental properties of these products, making this material approachable for a broad mathematical audience. Next, as one application of the cap product, we will show how it naturally arises in the analysis of automorphisms of groupoid $C^{*}$-algebras. As observed in [27, Section 5], a cocycle $\xi:\mathcal{G}\to\mathbb{T}$ (i.e. a homomorphism) gives rise to an automorphism $\alpha_{\xi}$ of the groupoid $C^{*}$-algebra $C^{*}_{r}(\mathcal{G})$, and $\alpha_{\xi}$ is inner if and only if the cohomology class of $\xi$ is zero in $H^{1}(\mathcal{G},\mathbb{T})$. In many concrete settings, the homomorphism $\xi:\mathcal{G}\to\mathbb{T}$ often lifts to a homomorphism $\mathcal{G}\to\mathbb{R}$, for example when $\mathcal{G}$ arises an action of $\mathbb{Z}^{N}$. In such a case, $\xi$ is homotopic to the constant map within cocycles, which implies that $\alpha_{\xi}$ is homotopic to the identity in $\operatorname{Aut}(C^{*}_{r}(\mathcal{G}))$. Then, it is natural to ask whether the automorphism $\alpha_{\xi}$ is asymptotically inner. Using the cap product, we introduce an invariant for $\xi$ which determines asymptotic innerness of $\alpha_{\xi}$ (Proposition 4.2 and Theorem 4.3).
@@ 2.1 Ample groupoids
In this article, by an étale groupoid we mean a second countable locally compact Hausdorff groupoid such that the range map is a local homeomorphism. (We emphasize that our étale groupoids are always assumed to be Hausdorff, while non-Hausdorff groupoids are also studied actively.) We refer the reader to [27, 28] for background material on étale groupoids. For an étale groupoid $\mathcal{G}$, we let $\mathcal{G}^{(0)}$ denote the unit space and let $s$ and $r$ denote the source and range maps, i.e. $s(g)=g^{-1}g$, $r(g)=gg^{-1}$. A subset $U\subset\mathcal{G}$ is called a bisection if $r|U$ and $s|U$ are injective. The étale groupoid $\mathcal{G}$ has a topology with a basis consisting of compact open bisections. For $x\in\mathcal{G}^{(0)}$, the set $r(s^{-1}(x))$ is called the orbit of $x$. When every orbit is dense in $\mathcal{G}^{(0)}$, $\mathcal{G}$ is said to be minimal. For a subset $Y\subset\mathcal{G}^{(0)}$, the reduction of $\mathcal{G}$ to $Y$ is $r^{-1}(Y)\cap s^{-1}(Y)$ and denoted by $\mathcal{G}|Y$. If $Y$ is open, then the reduction $\mathcal{G}|Y$ is an étale subgroupoid of $\mathcal{G}$ in an obvious way.
@@ 2.2 Homology groups and cohomology groups
Definition 2.1 ([5, Section 3.1],[18, Definition 3.1]).
@@ 3 Cup and cap products
In this section, we introduce the cup and cap products for ample groupoids. We start by introducing the cup product. Let $A$ be a topological abelian group.
@@
The following relationship holds between the cup and cap products.
@@ 4 Automorphisms of $C^{*}$-algebras
for $a\in A$ and $g\in\mathcal{G}$. Clearly $\alpha_{\xi}$ is inner if and only if $[\xi]=0$ in $H^{1}(\mathcal{G},\mathbb{T})$ (see [18, Proposition 5.7]). We examine conditions under which $\alpha_{\xi}$ is asymptotically inner.
@@
In the setting above, if $\alpha_{\xi}$ is asymptotically inner, then
@@
see [17] for instance. By [15, Theorem 4.3] and its proof, when $\alpha_{\xi}$ is asymptotically inner, we have $\Delta_{A}(u^{*}\alpha_{\xi}(u))=0$ for all $u\in U_{\infty}(A)$.
The groupoid $\mathcal{G}$ arises from a free action of $\mathbb{Z}^{N}$, and so $\mathcal{G}$ is almost finite by [18, Lemma 6.3]. Therefore, by [18, Theorem 7.5], any element $c$ of $H_{1}(\mathcal{G})$ is represented by a cycle of the form $1_{U}\in C_{c}(\mathcal{G},\mathbb{Z})$, where $U$ is a bisection such that $r(U)=\mathcal{G}^{(0)}=s(U)$. We may regard $1_{U}$ as a unitary in $A=C^{*}_{r}(\mathcal{G})$ By a straightforward computation, we see that $h:=(1_{U})^{*}\alpha_{\xi}(1_{U})$ is a continuous function on $\mathcal{G}^{(0)}$ such that $h(s(g))=\xi(g)$ for all $g\in U$. Pick $f\in C(\mathcal{G}^{(0)},\mathbb{R})$ such that $h(x)=e^{2\pi if(x)}$, and put $h_{t}(x):=e^{2\pi itf(x)}$ for $t\in[0,1]$. Then, $(h_{t})_{t\in[0,1]}$ is a smooth path of unitaries satisfying $h_{0}=1$ and $h_{1}=h$. As
@@
in $\operatorname{Aff}(M(\mathcal{G}))/\operatorname{Im}D_{\mathcal{G}}$. Since $\mathcal{G}$ is the transformation groupoid of a free minimal action of $\mathbb{Z}^{N}$, under the identification of $T(A)$ with $M(\mathcal{G})$, the gap labeling theorem [2, 3, 9] implies $\operatorname{Im}D_{A}=\operatorname{Im}D_{\mathcal{G}}$. Hence we obtain $\rho_{\mathcal{G}}\left([1_{U}]\frown[\xi]\right)=0$. ∎
@@
The implication (1$\implies$(2) is already shown in Proposition 4.2. We prove (2$\implies$(1). By [21, Corollary 5.7] (see also [12]), $A$ is a unital separable simple $C^{*}$-algebra with tracial rank zero and satisfying the UCT. We would like to apply Theorem 4.1 (see also [13, Theorem 3.1] and [15, Theorem 10.7]).
@@
It remains to show that the rotation map vanishes ([15, Section 3]). Consider the following short exact sequence of $C^{*}$-algebras:
@@
is the mapping torus of $\alpha_{\xi}$. We define a homomorphism $\beta:A\to M$ by $\beta(a)(t):=\alpha_{\xi_{t}}(a)$ for $t\in[0,1]$. Clearly $\beta$ is a right inverse of $\pi$. Let $R:K_{1}(M)\to\operatorname{Aff}(T(A))$ be the rotation map defined in [15, Definition 3.2]. By [15, Lemma 3.3], we have the following diagram:
@@
When $\mathcal{G}$ is the transformation groupoid of a minimal $\mathbb{Z}$-action, the theorem above has been shown in [16, Proposition 6.1]. In this case $H_{1}(\mathcal{G})$ is $\mathbb{Z}$.
@@ 5.3 The Penrose tiling
In this subsection, we compute a cup product of cohomology of the Penrose tiling. The argument relies on the results in [1]. Let $\Omega$ be the space of the Penrose tiling and let $\mathcal{G}$ be the associated ample groupoid. Our aim is to determine the cup product $H^{1}(\mathcal{G})\times H^{1}(\mathcal{G})\to H^{2}(\mathcal{G})$. There exists a two-dimensional CW-complex $\Gamma$ and a continuous surjection $\gamma:\Gamma\to\Gamma$ such that $\Omega$ is homeomorphic to the inverse limit of $(\Gamma,\gamma)$. As mentioned in [1, Section 10.4], $\gamma$ induces isomorphisms on $H^{i}(\Gamma)$, and so $H^{i}(\Omega)\cong H^{i}(\mathcal{G})$ is naturally isomorphic to $H^{i}(\Gamma)$. Thus, it suffices to compute the cup product $H^{1}(\Gamma)\times H^{1}(\Gamma)\to H^{2}(\Gamma)$.
@@
be the coboundary operators. Then, as computed in [1, Section 10.4], we get
@@ 5.4 The Ammann tiling
In the same way as the previous subsection, we compute the cup product $H^{1}\times H^{1}\to H^{2}$ for the Ammann tiling. Let $\Omega$ be the space of the Ammann tiling and let $\mathcal{G}$ be the associated ample groupoid. There exists a two-dimensional CW-complex $\Gamma$ and a continuous surjection $\gamma:\Gamma\to\Gamma$ such that $\Omega$ is homeomorphic to the inverse limit of $(\Gamma,\gamma)$. As mentioned in [1, Section 10.3], $\gamma$ induces isomorphisms on $H^{i}(\Gamma)$, and so $H^{i}(\Omega)\cong H^{i}(\mathcal{G})$ is naturally isomorphic to $H^{i}(\Gamma)$. Thus, it suffices to compute the cup product $H^{1}(\Gamma)\times H^{1}(\Gamma)\to H^{2}(\Gamma)$.
@@
and $V:=\{\alpha,\beta,\gamma\}$. The left-hand side of [1, Figure 4] shows its faces and how they are glued together. Let

### v3

@@ 1 Introduction
We aim to introduce the cup and cap products in a manner that is both thorough and accessible. Our goal is to ensure that even those who may not be familiar with algebraic topology can easily grasp these concepts. We will carefully explain the definitions and discuss the fundamental properties of these products, making this material approachable for a broad mathematical audience. Next, as one application of the cap product, we will show how it naturally arises in the analysis of automorphisms of groupoid $C^{*}$-algebras. As observed in [35, Section 5], a cocycle $\xi:\mathcal{G}\to\mathbb{T}$ (i.e. a homomorphism) gives rise to an automorphism $\alpha_{\xi}$ of the groupoid $C^{*}$-algebra $C^{*}_{r}(\mathcal{G})$, and $\alpha_{\xi}$ is inner if and only if the cohomology class of $\xi$ is zero in $H^{1}(\mathcal{G},\mathbb{T})$. In many concrete settings, the homomorphism $\xi:\mathcal{G}\to\mathbb{T}$ often lifts to a homomorphism $\mathcal{G}\to\mathbb{R}$, for example when $\mathcal{G}$ arises an action of $\mathbb{Z}^{N}$. In such a case, $\xi$ is homotopic to the constant map within cocycles, which implies that $\alpha_{\xi}$ is homotopic to the identity in $\operatorname{Aut}(C^{*}_{r}(\mathcal{G}))$. Then, it is natural to ask whether the automorphism $\alpha_{\xi}$ is asymptotically inner. Using the cap product, we introduce an invariant for $\xi$ which determines asymptotic innerness of $\alpha_{\xi}$ (Proposition 4.2 and Theorem 4.5).
@@
Acknowledgements.
We would like to thank the anonymous referee for many helpful comments, including pointing out prior work by Crainic and Moerdijk on the cohomology and homology of étale groupoids, as well as known issues concerning the proof of the gap labelling conjecture. We are also grateful to A. Miller and V. Proietti for kindly informing us about the current status of the gap labelling conjecture.
@@ 2.1 Ample groupoids
In this article, by an étale groupoid we mean a second countable locally compact Hausdorff groupoid such that the range map is a local homeomorphism. (We emphasize that our étale groupoids are always assumed to be Hausdorff, while non-Hausdorff groupoids are also studied actively.) We refer the reader to [35, 36] for background material on étale groupoids. For an étale groupoid $\mathcal{G}$, we let $\mathcal{G}^{(0)}$ denote the unit space and let $s$ and $r$ denote the source and range maps, i.e. $s(g)=g^{-1}g$, $r(g)=gg^{-1}$. A subset $U\subset\mathcal{G}$ is called a bisection if $r|U$ and $s|U$ are injective. The étale groupoid $\mathcal{G}$ has a topology with a basis consisting of open bisections. For $x\in\mathcal{G}^{(0)}$, the set $r(s^{-1}(x))$ is called the orbit of $x$. When every orbit is dense in $\mathcal{G}^{(0)}$, $\mathcal{G}$ is said to be minimal. For a subset $Y\subset\mathcal{G}^{(0)}$, the reduction of $\mathcal{G}$ to $Y$ is $r^{-1}(Y)\cap s^{-1}(Y)$ and denoted by $\mathcal{G}|Y$. If $Y$ is open, then the reduction $\mathcal{G}|Y$ is an étale subgroupoid of $\mathcal{G}$ in an obvious way.
@@ 2.2 Homology groups and cohomology groups
Definition 2.1 ([9, Section 3.1],[25, Definition 3.1]).
@@
In [9], for any étale groupoid $\mathcal{G}$ and any $\mathcal{G}$-sheaf $A$, the homology groups $H_{n}(\mathcal{G};A)$ are introduced. In this article, we restrict our attention to the case of constant sheaves.
@@
In [14], for any topological groupoid $\mathcal{G}$ and any $\mathcal{G}$-sheaf $A$, the cohomology groups $H^{n}(\mathcal{G};A)$ are introduced. In this article, we restrict our attention to the case of constant sheaves on étale groupoids.
@@ 3 Cup and cap products
In this section, we introduce the cup and cap products for ample groupoids. The cup and cap products have already been studied by Crainic and Moerdijk [8, 7] in the context of sheaves on étale groupoids (see Remarks 3.3 and 3.6 below). Our aim here is to provide a concrete and explicit model for these products in the setting of constant sheaves on ample groupoids, designed to be accessible without requiring extensive background in algebraic topology.
We start by introducing the cup product. Let $A$ be a topological abelian group.
@@
The cup product for étale groupoids has been studied in the literature via sheaf-theoretic approaches. To address this and provide clarity, we include the following remark.
Remark 3.3.
For any étale groupoid $\mathcal{G}$ and $\mathcal{G}$-sheaf $A$, Moerdijk [28] proved that Haefliger’s cohomology $H^{*}(\mathcal{G};A)$ is isomorphic to the sheaf cohomology $H^{*}(B\mathcal{G};\tilde{A})$, where $B\mathcal{G}$ is the classifying space of $\mathcal{G}$ and $\tilde{A}$ is the sheaf on $B\mathcal{G}$ induced by $A$. The sheaf cohomology has cup products (see [13, Section II.6.6] for instance), and the cup product introduced in the theorem above is its particular case. In [28], the isomorphism between $H^{*}(\mathcal{G};A)$ and $H^{*}(B\mathcal{G};\tilde{A})$ was constructed via the cohomology of the topological category $\operatorname{simp}(\mathcal{G})$, which is the topological sum of all $\mathcal{G}^{(n)}$, $n\geq 0$. When $A$ is a constant sheaf on an ample groupoid $\mathcal{G}$, one can check that the isomorphisms intertwine cup products. Another description of cup products is given in [8, Section 2.10] (which is the preprint version of [9]).
@@
As with the cup product, the cap product has a well-established abstract formulation. We therefore include the following remark to connect our concrete construction with previous works.
Remark 3.6.
A model of cap product has been constructed in Crainic’s PhD thesis ([7, Section 2.2.12]) in the following way. Let $\mathcal{G}$ be an étale groupoid. For any $\mathcal{G}$-sheaves $A,B$, one can construct $\operatorname{Ext}_{\mathcal{G}}^{*}(A,B)$ as the right derived functor of $B\mapsto\operatorname{Hom}_{\mathcal{G}}(A,B)$. The particular case is $\operatorname{Ext}_{\mathcal{G}}^{*}(\mathbb{Z},B)=H^{*}(\mathcal{G};B)$. As observed in [7, Section 1.2.6], the Ext groups act on the the homology:
$H_{n}(\mathcal{G};A)\times\operatorname{Ext}_{\mathcal{G}}^{m}(A,B)\to H_{n-m}(\mathcal{G};B).$
When working over $\mathbb{Z}$, this yields the cap product for étale groupoids:
$\frown\ :H_{n}(\mathcal{G};\mathbb{Z})\times H^{m}(\mathcal{G};A)\to H_{n-m}(\mathcal{G};A).$
Theorem 3.5 gives a concrete description of the cap product in the case of constant sheaves on ample groupoids.
The following relationship holds between the cup and cap products. We remark that this also follows from the Crainic’s presentation of cap products mentioned in the remark above. See [7, Section 1.2.6, 2.2.12].
@@ 4 Automorphisms of $C^{*}$-algebras
for $a\in A$ and $g\in\mathcal{G}$. Clearly $\alpha_{\xi}$ is inner if and only if $[\xi]=0$ in $H^{1}(\mathcal{G},\mathbb{T})$ (see [25, Proposition 5.7]). We examine conditions under which $\alpha_{\xi}$ is asymptotically inner.
@@
Let $N\leq 3$. If $\alpha_{\xi}$ is asymptotically inner, then
@@
see [24] for instance. By [22, Theorem 4.3] and its proof, when $\alpha_{\xi}$ is asymptotically inner, we have $\Delta_{A}(u^{*}\alpha_{\xi}(u))=0$ for all $u\in U_{\infty}(A)$.
The groupoid $\mathcal{G}$ arises from a free action of $\mathbb{Z}^{N}$, and so $\mathcal{G}$ is almost finite by [25, Lemma 6.3]. Therefore, by [25, Theorem 7.5], any element $c$ of $H_{1}(\mathcal{G})$ is represented by a cycle of the form $1_{U}\in C_{c}(\mathcal{G},\mathbb{Z})$, where $U$ is a bisection such that $r(U)=\mathcal{G}^{(0)}=s(U)$. We may regard $1_{U}$ as a unitary in $A=C^{*}_{r}(\mathcal{G})$. By a straightforward computation, we see that $h:=(1_{U})^{*}\alpha_{\xi}(1_{U})$ is a continuous function on $\mathcal{G}^{(0)}$ such that $h(s(g))=\xi(g)$ for all $g\in U$. Pick $f\in C(\mathcal{G}^{(0)},\mathbb{R})$ such that $h(x)=e^{2\pi if(x)}$, and put $h_{t}(x):=e^{2\pi itf(x)}$ for $t\in[0,1]$. Then, $(h_{t})_{t\in[0,1]}$ is a smooth path of unitaries satisfying $h_{0}=1$ and $h_{1}=h$. As
@@
in $\operatorname{Aff}(M(\mathcal{G}))/\operatorname{Im}D_{\mathcal{G}}$. Since $\mathcal{G}$ is the transformation groupoid of a free minimal action of $\mathbb{Z}^{N}$ with $N\leq 3$, under the identification of $T(A)$ with $M(\mathcal{G})$, the gap labeling theorem [4, 1] (see also the remark below) implies $\operatorname{Im}D_{A}=\operatorname{Im}D_{\mathcal{G}}$. Hence we obtain $\rho_{\mathcal{G}}\left([1_{U}]\frown[\xi]\right)=0$. ∎
Remark 4.3.
Let $\mathcal{G}$ be an ample groupoid arising from a free action of $\mathbb{Z}^{N}$ on the Cantor set and let $A:=C^{*}_{r}(\mathcal{G})$. The gap labelling conjecture asks whether or not the image of the map
$D_{\mathcal{G}}:H_{0}(\mathcal{G})\to\operatorname{Aff}(M(\mathcal{G}))$
and the image of the map
$D_{A}:K_{0}(A)\to\operatorname{Aff}(T(A))$
agree under the identification of $T(A)$ with $M(\mathcal{G})$. Several proofs for this conjecture were proposed ([3, 5, 16]), but it is known that they have issues (see page 1663 of [12] for instance). This is why we need to assume $N\leq 3$ in the proposition above.
Remark 4.4.
Given a cocycle $\xi\in C(\mathcal{G},\mathbb{T})$, one can consider a $\mathcal{G}$-sheaf $A$ of germs of continuous ×$\mathbb{Z}\times\mathbb{T}$-valued functions such that the $\mathcal{G}$-action is given by the translation by $\xi$ on the $\mathbb{T}$-component. More precisely, for any compact open bisection $U\subset\mathcal{G}$, the $\mathcal{G}$-action is described as follows:
$\left((f_{1},f_{2})\cdot g\right)(r(g))=\left(f_{1}(s(g)),f_{2}(s(g))+f_{1}(s(g))\xi(g)\right)$
for $f_{1}\in C(s(U),\mathbb{Z})$, $f_{2}\in C(s(U),\mathbb{T})$ and $g\in U$. By [9, Section 3.6], we obtain the long exact sequence in homology:
[…]
The cap product $c\mapsto c\frown[\xi]$ appeared in the proof of Proposition 4.2 equals the homomorphism $H_{1}(\mathcal{G},\mathbb{Z})\to H_{0}(\mathcal{G},\mathbb{T})$ in this long exact sequence.
@@
The implication (1$\implies$(2) is already shown in Proposition 4.2. We prove (2$\implies$(1). By [29, Theorem B] (see also [19]), $A$ is a unital separable simple $C^{*}$-algebra with tracial rank zero and satisfying the UCT. We would like to apply Theorem 4.1 (see also [20, Theorem 3.1] and [22, Theorem 10.7]).
@@
It remains to show that the rotation map vanishes ([22, Section 3]). Consider the following short exact sequence of $C^{*}$-algebras:
@@
is the mapping torus of $\alpha_{\xi}$. We define a homomorphism $\beta:A\to M$ by $\beta(a)(t):=\alpha_{\xi_{t}}(a)$ for $t\in[0,1]$. Clearly $\beta$ is a right inverse of $\pi$. Let $R:K_{1}(M)\to\operatorname{Aff}(T(A))$ be the rotation map defined in [22, Definition 3.2]. By [22, Lemma 3.3], we have the following diagram:
@@
When $\mathcal{G}$ is the transformation groupoid of a minimal $\mathbb{Z}$-action, the theorem above has been shown in [23, Proposition 6.1]. In this case $H_{1}(\mathcal{G})$ is $\mathbb{Z}$.
@@ 5.3 The Penrose tiling
In this subsection, we compute a cup product of cohomology of the Penrose tiling. The argument relies on the results in [2]. Let $\Omega$ be the space of the Penrose tiling and let $\mathcal{G}$ be the associated ample groupoid. Our aim is to determine the cup product $H^{1}(\mathcal{G})\times H^{1}(\mathcal{G})\to H^{2}(\mathcal{G})$. There exists a two-dimensional CW-complex $\Gamma$ and a continuous surjection $\gamma:\Gamma\to\Gamma$ such that $\Omega$ is homeomorphic to the inverse limit of $(\Gamma,\gamma)$. As mentioned in [2, Section 10.4], $\gamma$ induces isomorphisms on $H^{i}(\Gamma)$, and so $H^{i}(\Omega)\cong H^{i}(\mathcal{G})$ is naturally isomorphic to $H^{i}(\Gamma)$. Thus, it suffices to compute the cup product $H^{1}(\Gamma)\times H^{1}(\Gamma)\to H^{2}(\Gamma)$.
@@
be the coboundary operators. Then, as computed in [2, Section 10.4], we get
@@ 5.4 The Ammann tiling
In the same way as the previous subsection, we compute the cup product $H^{1}\times H^{1}\to H^{2}$ for the Ammann tiling. Let $\Omega$ be the space of the Ammann tiling and let $\mathcal{G}$ be the associated ample groupoid. There exists a two-dimensional CW-complex $\Gamma$ and a continuous surjection $\gamma:\Gamma\to\Gamma$ such that $\Omega$ is homeomorphic to the inverse limit of $(\Gamma,\gamma)$. As mentioned in [2, Section 10.3], $\gamma$ induces isomorphisms on $H^{i}(\Gamma)$, and so $H^{i}(\Omega)\cong H^{i}(\mathcal{G})$ is naturally isomorphic to $H^{i}(\Gamma)$. Thus, it suffices to compute the cup product $H^{1}(\Gamma)\times H^{1}(\Gamma)\to H^{2}(\Gamma)$.
@@
and $V:=\{\alpha,\beta,\gamma\}$. The left-hand side of [2, Figure 4] shows its faces and how they are glued together. Let